1. what is the equivalent resistance of this circuit?

(250 ohms, 150 ohms, and 50 ohms are their in parallel)


1
Expert's answer
2021-05-27T10:18:14-0400

Given,

"R_1=250\ \Omega\\R_2=150\ \Omega\\R_3=50\ \Omega"


and all the three resistors are in parallel

So, Equivalent Resistance

"\dfrac{1}{R_{eq}}=\dfrac{1}{R_1}+\dfrac{1}{R_2}+\dfrac{1}{R_3}\\\ \\\dfrac{1}{R_{eq}}=\dfrac{1}{250}+\dfrac{1}{150}+\dfrac{1}{50}\\\ \\\dfrac{1}{R_{eq}}=\dfrac{3+5+15}{750}\\\ \\R_{eq}= \dfrac{750}{23}=32.61"


Hence, equivalent resistance "\boxed{R_{eq}=32.61\ \Omega}"